one canned juice drink 30% orange juice, another is 10% orange juice. How many liters of each should be mixed togethr in order to get 20L that is 19% orange juice.
How many liters of the 10% orange juice should be in the mixture?

Answer by tutorcecilia(2152)

x + y = 20
So, y = 20-x
And,
.30x + .10y = (20)(.19)
.30x + .10 (20-x) = 3.8
.30x + 2 - .10x = 3.8
.30x - .10x = 3.8 - 2
.20x=1.8
x=1.8/.20
x=9 liters at 30%
So,
x+y=20
9+y=20
y=20-9
y=11 liters at 10%